9:50 PM Wed Jan 30 ES316 assignments 1.pdf Engineering Economics 4 The Learning Curve The Mountain M bike workshop can only sell exact ly one customized bike every month, becausethe city only agrees to issue one plate for this type of modiied bike per mouth. So they make exact ly one every moath. Dg hefirst year of business (12 bikes was made), on average they can make a bike in 100 labor urs During the second year of business (anot her 12 was made), they improved a lot and were able to on average make the same bike in 75 labor hours. Unfort unately, we do not have records on how many labor hours they spent to make the bike in each month of the 24 months. Please estimate the learning-curve percent age/rate for their bike production. You may make any assumpt lons as you need, but please just ify them so that they sound reasonable to other engineers in the room CS) Dashboard in calendar -ToDo Notifications Inbox
Answer #1

In this case, we do not have the monthly hours spent on each bike and have only annual averages. So, we can consider them as being produced in the batches. Each batch has 12 bikes.

We observe that average production time has decreased in this case. To explain the process, we assume that

1.Time required to complete a job or task will be less each time same job/task is taken. Each individual or even organization learns by practice. More the repetitive jobs are done, more efficiently people work. They take less time in future for doing the same job.In this case time required for each batch of bikes (average time for each bike production) will decrease as new batches are produced.

2.Time required per batch will decrease but at decreasing rate. In the second year of production, average time has decreased by 25 hours.In third batch average production time will not decrease by 25 hours but may be by 18 hours or 20 hours.

3.Reduction in time follows a pattern.

Time taken for first batch in year 1=B1=100*12=1200 hours

Time taken for second batch in year 2=B2=75*12=900 hour/bike

Learning Rate=B2//B1=900/1200=0.75 or 75%